# Paired kernels and truncated Toeplitz operators

### v1

@@ 4 Kernels of ATTO
$\phi=\overline{\theta}\frac{A_{1-}}{B_{1-}}-\alpha\frac{A_{2+}}{B_{2+}}\in L^{\infty},$ (4.1)
@@
$\mathcal{K}=(\tilde{f}^{T}\mathop{\rm ker}\nolimits H_{f\tilde{f}^{T}})\cap(\tilde{g}^{T}\mathop{\rm ker}\nolimits\widetilde{H}_{g\tilde{g}^{T}}),$ (4.7)
@@
$\mathop{\rm ker}\nolimits H_{f\tilde{f}^{T}}=\{\phi_{+}\in(H^{2}_{+})^{2}:P^{-}f\tilde{f}^{T}\phi_{+}=0\}$ (4.8)
@@
$\mathop{\rm ker}\nolimits\widetilde{H}_{g\tilde{g}^{T}}=\{\phi_{-}\in(H^{2}_{-})^{2}:P^{+}g\tilde{g}^{T}\phi_{-}=0\}.$ (4.9)
@@
$\mathop{\rm ker}\nolimits T_{G}=\mathcal{K}f.$ (4.10)
@@
$G=\begin{bmatrix}\overline{\theta}&0\\ \phi&\alpha\end{bmatrix}.$ (4.11)
@@
$\mathop{\rm ker}\nolimits A^{\theta,\alpha}_{\phi}=P_{1}\mathop{\rm ker}\nolimits T_{G}.$ (4.12)
@@
$f=\theta\begin{bmatrix}1\\ \frac{A_{2+}}{B_{2+}}\end{bmatrix},\qquad g=\begin{bmatrix}1\\ \frac{A_{1-}}{B_{1-}}\end{bmatrix},$ (4.13)
@@
$\tilde{f}=\begin{bmatrix}\,\overline{\theta}\,\\ 0\end{bmatrix},\qquad\tilde{g}=\begin{bmatrix}1\\ 0\end{bmatrix}.$ (4.14)
@@
$G=\begin{bmatrix}\overline{\theta}&0\\ \overline{\theta}\frac{A_{1-}}{B_{1-}}-\alpha\frac{A_{2+}}{B_{2+}}&\alpha\end{bmatrix}$ (4.15)
@@
$\mathcal{S}:=\left(\alpha\begin{bmatrix}1&\frac{A_{2+}}{B_{2+}}\end{bmatrix}(H^{2}_{+})^{2}\right)\cap\left(\begin{bmatrix}1&\frac{A_{1-}}{B_{1-}}\end{bmatrix}(H^{2}_{-})^{2}\right)=\{0\},$ (4.16)
@@
$\mathop{\rm ker}\nolimits A^{\theta,\alpha}_{\phi}=\mathcal{K}\theta,$ (4.17)
@@
$\mathcal{K}=\biggl{(}\begin{bmatrix}\,\overline{\theta}&0\end{bmatrix}\underbrace{\mathop{\rm ker}\nolimits H_{\begin{bmatrix}1&0\\ \frac{A_{2+}}{B_{2+}}&0\end{bmatrix}}}_{\mathop{\rm ker}\nolimits H_{f\tilde{f}^{T}}}\biggr{)}\cap\biggl{(}\begin{bmatrix}1&0\end{bmatrix}\underbrace{\mathop{\rm ker}\nolimits\widetilde{H}_{\begin{bmatrix}1&0\\ \frac{A_{1-}}{B_{1-}}&0\end{bmatrix}}}_{\mathop{\rm ker}\nolimits\widetilde{H}_{g\tilde{g}^{T}}}\biggr{)}.$ (4.18)
@@
$B_{2+}\mathcal{S}\subseteq\alpha{\rm ker}^{+}S_{\alpha B_{1-},B_{2+}}.$ (4.19)
@@
$\alpha\left(\phi_{1+}+\frac{A_{2+}}{B_{2+}}\phi_{2+}\right)\qquad\hbox{with}\quad\phi_{1+},\phi_{2+}\in H^{2}_{+},$ (4.20)
$\alpha\left(\phi_{1+}+\frac{A_{2+}}{B_{2+}}\phi_{2+}\right)=\phi_{1-}+\frac{A_{1-}}{B_{1-}}\phi_{2-}\qquad\hbox{for some}\quad\phi_{1-},\phi_{2-}\in H^{2}_{-},$ (4.21)
@@
$B_{1-}B_{2+}\left(\alpha(\phi_{1+}+\frac{A_{2+}}{B_{2+}}\phi_{2+})\right)$ $=$ $(\alpha B_{1-}(B_{2+}\phi_{1+}+A_{2+}\phi_{2+}))$
@@
$B_{1-}B_{2+}\mathcal{S}\subseteq\alpha B_{1-}{\rm ker}^{+}S_{\alpha B_{1-},B_{2+}}.$
@@
$\mathcal{S}=\{0\}\quad\hbox{if}\quad\mathop{\rm ker}\nolimits T_{\alpha\overline{I}}=\{0\},$ (4.22)
@@
$\mathcal{H}_{X}=\{h\in\mathcal{H}:Xh\in\mathcal{H}\}.$ (4.24)
@@
$(H^{2}_{-})_{A_{1-}/B_{1-}}=\mathop{\rm ker}\nolimits\widetilde{H}_{A_{1-}/B_{1-}}=B_{1-}H^{2}_{-}.$ (4.26)
@@
$\widetilde{A}_{2+}A_{2+}\phi_{+}=\widetilde{A}_{2+}B_{2+}\psi_{+}$ $\iff$ $\phi_{1+}-B_{2+}\widetilde{B}_{2_{+}}\phi_{1+}=\widetilde{A}_{2+}B_{2+}\psi_{+}$
@@
$\mathop{\rm ker}\nolimits H_{f\tilde{f}^{T}}=(B_{2+}H^{2}_{+},H^{2}_{+})$ (4.27)
@@
$\mathop{\rm ker}\nolimits\widetilde{H}_{g\tilde{g}^{T}}=(B_{1-}H^{2}_{-},H^{2}_{-}).$ (4.28)
@@
$\mathop{\rm ker}\nolimits A^{\theta,\alpha}_{\phi}=B_{2+}{\rm ker}^{+}S_{\overline{\theta}B_{2+},-B_{1-}}.$ (4.29)
@@
$\mathop{\rm ker}\nolimits A^{\theta,\alpha}_{\phi}=I_{2}\mathop{\rm ker}\nolimits T_{\overline{\theta}(I_{1}I_{2})}.$ (4.30)
@@
$\phi=\overline{\theta}R_{+}-\alpha R_{-}+\sum_{j=1}^{N}\frac{\overline{\theta}P^{\alpha}_{n_{j}-1}(z)-\alpha P^{\overline{\theta}}_{n_{j}-1}(t_{j})}{(z-t_{j})^{n_{j}}},$ (4.31)
@@
$\phi=\overline{\theta}\frac{Q_{1}}{\mathcal{E}D_{1+}}-\alpha\frac{Q_{2}}{\mathcal{E}D_{2-}}\in L^{\infty},$ (4.32)
@@
$\phi=\overline{\theta}\frac{A_{1-}}{B_{1-}}-\alpha\frac{A_{2+}}{B_{2+}},$ (4.33)

### v2

@@ 4. Kernels of ATTO
(4.1) $\phi=\overline{\theta}\frac{A_{1-}}{B_{1-}}-\alpha\frac{A_{2+}}{B_{2+}}\in L^{\infty},$
@@
(4.7) $\mathcal{K}=(\tilde{f}^{T}\mathop{\rm ker}\nolimits H_{f\tilde{f}^{T}})\cap(\tilde{g}^{T}\mathop{\rm ker}\nolimits\widetilde{H}_{g\tilde{g}^{T}}),$
@@
(4.8) $\mathop{\rm ker}\nolimits H_{f\tilde{f}^{T}}=\{\phi_{+}\in(H^{2}_{+})^{2}:P^{-}f\tilde{f}^{T}\phi_{+}=0\}$
@@
(4.9) $\mathop{\rm ker}\nolimits\widetilde{H}_{g\tilde{g}^{T}}=\{\phi_{-}\in(H^{2}_{-})^{2}:P^{+}g\tilde{g}^{T}\phi_{-}=0\}.$
@@
(4.10) $\mathop{\rm ker}\nolimits T_{G}=\mathcal{K}f.$
@@
(4.11) $G=\begin{bmatrix}\overline{\theta}&0\\ \phi&\alpha\end{bmatrix}.$
@@
(4.12) $\mathop{\rm ker}\nolimits A^{\theta,\alpha}_{\phi}=P_{1}\mathop{\rm ker}\nolimits T_{G}.$
@@
(4.13) $f=\theta\begin{bmatrix}1\\ \frac{A_{2+}}{B_{2+}}\end{bmatrix},\qquad g=\begin{bmatrix}1\\ \frac{A_{1-}}{B_{1-}}\end{bmatrix},$
@@
(4.14) $\tilde{f}=\begin{bmatrix}\,\overline{\theta}\,\\ 0\end{bmatrix},\qquad\tilde{g}=\begin{bmatrix}1\\ 0\end{bmatrix}.$
@@
(4.15) $G=\begin{bmatrix}\overline{\theta}&0\\ \overline{\theta}\frac{A_{1-}}{B_{1-}}-\alpha\frac{A_{2+}}{B_{2+}}&\alpha\end{bmatrix}$
@@
(4.16) $\mathcal{S}:=\left(\alpha\begin{bmatrix}1&\frac{A_{2+}}{B_{2+}}\end{bmatrix}(H^{2}_{+})^{2}\right)\cap\left(\begin{bmatrix}1&\frac{A_{1-}}{B_{1-}}\end{bmatrix}(H^{2}_{-})^{2}\right)=\{0\},$
@@
(4.17) $\mathop{\rm ker}\nolimits A^{\theta,\alpha}_{\phi}=\mathcal{K}\theta,$
@@
(4.18) $\mathcal{K}=\biggl{(}\begin{bmatrix}\,\overline{\theta}&0\end{bmatrix}\underbrace{\mathop{\rm ker}\nolimits H_{\begin{bmatrix}1&0\\ \frac{A_{2+}}{B_{2+}}&0\end{bmatrix}}}_{\mathop{\rm ker}\nolimits H_{f\tilde{f}^{T}}}\biggr{)}\cap\biggl{(}\begin{bmatrix}1&0\end{bmatrix}\underbrace{\mathop{\rm ker}\nolimits\widetilde{H}_{\begin{bmatrix}1&0\\ \frac{A_{1-}}{B_{1-}}&0\end{bmatrix}}}_{\mathop{\rm ker}\nolimits\widetilde{H}_{g\tilde{g}^{T}}}\biggr{)}.$
@@
(4.19) $B_{2+}\mathcal{S}\subseteq\alpha{\rm ker}^{+}S_{\alpha B_{1-},B_{2+}}.$
@@
(4.20) $\alpha\left(\phi_{1+}+\frac{A_{2+}}{B_{2+}}\phi_{2+}\right)\qquad\hbox{with}\quad\phi_{1+},\phi_{2+}\in H^{2}_{+},$
(4.21) $\alpha\left(\phi_{1+}+\frac{A_{2+}}{B_{2+}}\phi_{2+}\right)=\phi_{1-}+\frac{A_{1-}}{B_{1-}}\phi_{2-}\qquad\hbox{for some}\quad\phi_{1-},\phi_{2-}\in H^{2}_{-},$
@@
$B_{1-}B_{2+}\left(\alpha(\phi_{1+}+\frac{A_{2+}}{B_{2+}}\phi_{2+})\right)$ $=\alpha B_{1-}(B_{2+}\phi_{1+}+A_{2+}\phi_{2+})$
@@
$B_{1-}B_{2+}\mathcal{S}\subseteq\alpha B_{1-}{\rm ker}^{+}S_{\alpha B_{1-},B_{2+}}$
as required. ∎
@@
(4.22) $\mathcal{S}=\{0\}\quad\hbox{if}\quad\mathop{\rm ker}\nolimits T_{\alpha\overline{I}}=\{0\},$
@@
(4.24) $\mathcal{H}_{X}=\{h\in\mathcal{H}:Xh\in\mathcal{H}\}.$
@@
(4.26) $(H^{2}_{-})_{A_{1-}/B_{1-}}=\mathop{\rm ker}\nolimits\widetilde{H}_{A_{1-}/B_{1-}}=B_{1-}H^{2}_{-}.$
@@
$\widetilde{A}_{2+}A_{2+}\phi_{1+}=\widetilde{A}_{2+}B_{2+}\psi_{+}$ $\iff\phi_{1+}-B_{2+}\widetilde{B}_{2_{+}}\phi_{1+}=\widetilde{A}_{2+}B_{2+}\psi_{+}$
@@
(4.27) $\mathop{\rm ker}\nolimits H_{f\tilde{f}^{T}}=(B_{2+}H^{2}_{+},H^{2}_{+})$
@@
(4.28) $\mathop{\rm ker}\nolimits\widetilde{H}_{g\tilde{g}^{T}}=(B_{1-}H^{2}_{-},H^{2}_{-}).$
@@
(4.29) $\mathop{\rm ker}\nolimits A^{\theta,\alpha}_{\phi}=B_{2+}{\rm ker}^{+}S_{\overline{\theta}B_{2+},-B_{1-}}.$
@@
(4.30) $\mathop{\rm ker}\nolimits A^{\theta,\alpha}_{\phi}=I_{2}\mathop{\rm ker}\nolimits T_{\overline{\theta}(I_{1}I_{2})}.$
@@
(4.31) $\phi=\overline{\theta}R_{+}-\alpha R_{-}+\sum_{j=1}^{N}\frac{\overline{\theta}P^{\alpha}_{n_{j}-1}(z)-\alpha P^{\overline{\theta}}_{n_{j}-1}(t_{j})}{(z-t_{j})^{n_{j}}},$
@@
(4.32) $\phi=\overline{\theta}\frac{Q_{1}}{\mathcal{E}D_{1+}}-\alpha\frac{Q_{2}}{\mathcal{E}D_{2-}}\in L^{\infty},$
@@
(4.33) $\phi=\overline{\theta}\frac{A_{1-}}{B_{1-}}-\alpha\frac{A_{2+}}{B_{2+}},$
@@
Acknowledgment
Work partially funded by FCT/Portugal through project UIDB/04459/2020 with DOI identifier 10-54499/UIDP/04459/2020